OpenStudy (countrygirl15):

Please help me solve for x and y! 2x+3y=4 5x+8y=11

OpenStudy (countrygirl15):

@amistre64 could you help me?

OpenStudy (campbell_st):

multiply the 1st equation by 5 and the 2nd equation by - 2 10x + 15y = 20 -10x -16y = -22 --------------------- Adding the equations will eliminate x and you can solve for y. when you have a value for y, substitute into either of the original equations to find x

OpenStudy (countrygirl15):

is y -1?

OpenStudy (campbell_st):

well after adding the equations you get -y = -2 so divide both by -1 then y = 2 substitute this into either of the original equations to find x. hope it helps

OpenStudy (countrygirl15):

ok I got x=-1 and y=2

OpenStudy (campbell_st):

that's correct well done

OpenStudy (countrygirl15):

Thank you for your help! :) @campbell_st